40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 196     Entry time: Sun Jul 6 02:45:56 2014
Author: Koji 
Type: Optics 
Category: General 
Subject: FSR Measurement 

3rd OMC FSR / Finesse measurement

RF AM was injected by detuning a HWP.

Attachment 1: finesse_measurements_log.pdf  91 kB  Uploaded Sun Jul 6 04:41:16 2014  | Show | Show all
ELOG V3.1.3-